P invested Rs 8000 for 2 years at the rate of 8% simple interest per annum and Q invested the same amount for the same time at the rate of 10% annual compound interest then calculate,
I) interest received by P
II) interest received by Q
III) Without altering time and rate of interest,how much more or less money should P have to invest for equal interests?

Grade:10

1 Answers

Arun
4 years ago
8000(1+8/100) = 80 * 108 = 8640 + 640 = 9280
 
While Q have total amount after 2 years
= 8000(1+ 10/100)^2 = 8000* 121/100 = 9680
 
Hence interest received by P = 1280
Interest received by Q = 1680